# Quantum Enhancement of Thermalization

### v1

@@ Acknowledgements.
- Greiner et al. [2002] M. Greiner, O. Mandel, T. Esslinger, T. W. Hänsch, and I. Bloch, Quantum phase transition from a superfluid to a Mott insulator in a gas of ultracold atoms, nature 415, 39 (2002).
@@
- Iubini et al. [2019] S. Iubini, L. Chirondojan, G.-L. Oppo, A. Politi, and P. Politi, Dynamical freezing of relaxation to equilibrium, Physical review letters 122, 084102 (2019).
@@
- Qiao and Grossmann [2021] Y. Qiao and F. Grossmann, Exact variational dynamics of the multimode Bose-Hubbard model based on su (m) coherent states, Physical Review A 103, 042209 (2021).
@@
- Pitaevskii [1961] L. P. Pitaevskii, Vortex lines in an imperfect Bose gas, Sov. Phys. JETP 13, 451 (1961).
- Einstein [1917] A. Einstein, Zum Quantensatz von Sommerfeld und Epstein, Verh. d. D. Physik. Ges , 82 (1917).
- Brillouin [1926] L. Brillouin, Remarques sur la mécanique ondulatoire, J. Phys. Radium 7, 353 (1926).
- Keller [1958] J. B. Keller, Corrected Bohr-Sommerfeld quantum conditions for nonseparable systems, Annals of Physics 4, 180 (1958).
@@
- [51] See Supplemental Material, which includes references [63, 64, 17, 65], at [url will be inserted by publisher] for a comprehensive account of numerical and analytical details including convergence of quantum simulations, the calculation of thermal averages in the disordered rings, and comparisons performed in the superfluid regime, .

### v2

@@ Acknowledgements.
- Greiner et al. [2002a] M. Greiner, O. Mandel, T. Esslinger, T. W. Hänsch, and I. Bloch, Quantum phase transition from a superfluid to a Mott insulator in a gas of ultracold atoms, nature 415, 39 (2002a).
@@
- Richaud and Penna [2018] A. Richaud and V. Penna, Phase separation can be stronger than chaos, New Journal of Physics 20, 105008 (2018).
@@
- Iubini et al. [2019] S. Iubini, L. Chirondojan, G.-L. Oppo, A. Politi, and P. Politi, Dynamical freezing of relaxation to equilibrium, Phys. Rev. Lett. 122, 084102 (2019).
@@
- Qiao and Grossmann [2021] Y. Qiao and F. Grossmann, Exact variational dynamics of the multimode Bose-Hubbard model based on SU() coherent states, Physical Review A 103, 042209 (2021).
- Qiao [2024] Y. Qiao, Dynamical multi-configuration generalized coherent states approach to many-body bosonic quantum systems, PhD thesis, Technische Universität Dresden, Dresden, Germany (2024), available at https://nbn-resolving.org/urn:nbn:de:bsz:14-qucosa2-921049.
@@
- Steel et al. [1998] M. J. Steel, M. Olsen, L. Plimak, P. Drummond, S. Tan, M. Collett, D. Walls, and R. Graham, Dynamical quantum noise in trapped bose-einstein condensates, Physical Review A 58, 4824 (1998).
- Richter et al. [2022] K. Richter, J. D. Urbina, and S. Tomsovic, Semiclassical roots of universality in many-body quantum chaos, Journal of Physics A: Mathematical and Theoretical 55, 453001 (2022).
@@
- Löck et al. [2010] S. Löck, A. Bäcker, R. Ketzmerick, and P. Schlagheck, Regular-to-chaotic tunneling rates: From the quantum to the semiclassical regime, Phys. Rev. Lett. 104, 114101 (2010).
- Schlagheck et al. [2011] P. Schlagheck, A. Mouchet, and D. Ullmo, Resonance-assisted tunneling in mixed regular–chaotic systems, in Dynamical Tunneling: Theory and Experiment, edited by S. Keshavamurthy and P. Schlagheck (Taylor & Francis CRC, 2011).
@@
- Greiner et al. [2002b] M. Greiner, O. Mandel, T. W. Hänsch, and I. Bloch, Collapse and revival of the matter wave field of a Bose–Einstein condensate, Nature 419, 51 (2002b).
- [54] See Supplemental Material, which includes references [73, 74, 18, 75, 76], at [url will be inserted by publisher] for a comprehensive account of numerical and analytical details including convergence of quantum simulations, the calculation of thermal averages in the disordered rings, comparisons performed in the regime of weak interactions, and a study of transition probabilities between near resonant states that sheds light on the mechanism of QET.
@@
- Note [1] We have verified that the Lyapunov spectra and their associated Kolmogorov-Sinai entropies are quantitatively similar for several phase-space points sampled according to the Wigner function of the initial state.
@@
- Deutsch [1991] J. M. Deutsch, Quantum statistical mechanics in a closed system, Physical review a 43, 2046 (1991).
- Srednicki [1994] M. Srednicki, Chaos and quantum thermalization, Physical Review E 50, 888 (1994).
@@
- Note [2] Note that in the strong hopping regime this mechanism is expected to become ineffective as compared to transport via classical channels, in close analogy to chaos-assisted tunneling suppressing ordinary tunneling channels between symmetric wells [31].
- Bardet et al. [2023] I. Bardet, Á. Capel, L. Gao, A. Lucia, D. Pérez-García, and C. Rouzé, Rapid thermalization of spin chain commuting hamiltonians, Physical Review Letters 130, 060401 (2023).
- Kumar and Choudhury [2024] S. Kumar and S. Choudhury, Prethermalization in aperiodically driven classical spin systems, arXiv preprint arXiv:2404.10224 (2024).
- Roushan et al. [2017] P. Roushan, C. Neill, J. Tangpanitanon, V. M. Bastidas, A. Megrant, R. Barends, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, A. Fowler, B. Foxen, M. Giustina, E. Jeffrey, J. Kelly, E. Lucero, J. Mutus, M. Neeley, C. Quintana, D. Sank, A. Vainsencher, J. Wenner, T. White, H. Neven, D. G. Angelakis, and J. Martinis, Spectroscopic signatures of localization with interacting photons in superconducting qubits, Science 358, 1175 (2017).
@@
- Dell’Anna [2022] L. Dell’Anna, Entanglement properties and ground-state statistics of free bosons, Phys. Rev. A 105, 032412 (2022).